How many different ways can you arrange the fo?
llowing four digit number 3239
thank you

Answer by Edwin McCravy(20055)

If we could tell the difference between the 3's, like as if it were 

3239, the answer would be 4!, 

but since we can't tell the difference between the 3's, we must divide by
the number of ways the 3's can be swapped within each 4-digit number, which
is 2!.

Answer 4!/2! = (4×3×2×1)/(2×1) = 24/2 = 12 ways:

Here they are:

 1. 2339
 2. 2393
 3. 2933
 4. 3239
 5. 3293  
 6. 3329
 7. 3392
 8. 3923
 9. 3932
10. 9233
11. 9323
12. 9332

Edwin